# Interplay between local and global order controls bifurcations in nonreciprocal systems without a conservation law

### v2

@@ Abstract
Linear stability analysis.—We focus on the symmetric nonreciprocal setting $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, also known as the perfectly nonreciprocal [15] setting. We start with the linear stability analysis of Eqs. (5) around their steady state. The trivial steady state corresponds to the disordered state with $m^{\mu}_{s}=0$, $q^{ab}_{s}=0$, and $q^{\mu\mu}_{s}\equiv q_{s}(J,K)\neq 0$ which is explicitly given in [55]. Up to first order, small perturbations $\delta\mathbf{m}(t)\equiv(\delta m^{a}(t),\delta m^{b}(t))$ decouple from perturbations $\delta\mathbf{q}(t)\equiv(\delta q^{aa}(t),\delta q^{bb}(t),\delta q^{ab}(t))$, and we obtain the linear equation $\tau d\delta\mathbf{m}(t)/dt=\mathbf{M}(q_{s};J,K)\delta\mathbf{m}(t)$ (see [55] for the linear equation for $\delta\mathbf{q}(t)$). The elements of the linear stability matrix, $M_{\mu\nu}(q_{s};J,K)$, read
$\!\!M_{aa}$ $=M_{bb}=\frac{q_{s}/q_{\rm crit}-1}{1+q_{s}}[1-2\sum\nolimits_{l,n}\!\!\overline{\mathcal{P}}^{+}_{l}\tanh(U^{a}_{l,n})],$
$\!\!M_{ab}$ $={-}M_{ba}{=}\sum\nolimits_{l,n}\!(2n{-}1)[\overline{\mathcal{P}}^{+}_{l}{+}\overline{\mathcal{P}}^{-}_{l}]\tanh(U^{a}_{l,n}),$ (7)
@@
where $\overline{\mathcal{P}}^{\pm}_{l}(q_{s})\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}$ are the probabilities (6) evaluated at steady state values, and we introduced the critical local order $q_{\rm crit}\equiv 1/(z-1)$, which depends on the coordination number of the lattice. The solution of the linear stability equation can be written as $\mathbf{\delta m}(t)=\sum_{k=\pm}\mathcal{A}_{k}\mathrm{e}^{\lambda_{k}t/\tau}\boldsymbol{\nu}_{k}$, where $\mathcal{A}_{\pm}$ are set by the initial conditions, $\boldsymbol{\nu}_{\pm}$ are the eigenvectors of $\mathbf{M}(q_{s};J,K)$ and $\lambda_{\pm}(q_{s};J,K)$ the corresponding eigenvalues
$\lambda_{\pm}(q_{s};J,K)=M_{aa}(q_{s};J,K)\pm{\rm i}M_{ab}(q_{s};J,K),$ (8)
@@
$\rm i$ being the imaginary unit. Since ${M_{ab}(q_{s};J,K\neq 0)\neq 0}$ (see proof in [55]), the eigenvalues are complex for $K\neq 0$ (see dashed lines in Fig. 1b), resulting in oscillatory perturbations. The Hopf bifurcation [61], also called type-${\rm II}_{\rm o}$ instability [62], occurs when complex conjugate eigenvalues transit the imaginary axis in the complex plane. According to Eq. (8) this occurs when $M_{aa}(q_{s};J,K)=0$ implying ${q_{s}(J,K)=q_{\rm crit}}$ as seen from Eq. (7). The Hopf bifurcation is thus set by the critical value $q_{\rm crit}$ for local order, and for $q_{s}>q_{\rm crit}$ we have ${\rm Re}(\lambda_{\pm})>0$ (solid line in Fig. 1b). In other words, when spins on the respective lattices are sufficiently aligned, a transition to an oscillatory state occurs (Fig. 1c,d). The critical value of the local order that determines the onset of coherent oscillations is our second main result that generalizes to other approximation schemes beyond the mean field approximation (see Appendix B). After sufficient local order is attained within the lattices, the frustration due to the nonreciprocal coupling gives rise to coherent oscillations: for $K>0$ a spin $\sigma^{a}_{i}$ wants to align with $\sigma^{b}_{i}$ that in turn tends to misalign with $\sigma^{a}_{i}$. This dynamical frustration results in an oscillatory motion of the order parameters [15]. Notably, for the one-dimensional lattice ($z=2$) we find $q_{\rm crit}=1$, which, in contrast to the mean-field prediction [48] (see also Appendix A), correctly implies the non-existence of a Hopf bifurcation.
@@
Figure 2: (a,b) Bifurcation diagram of the magnetization $(m^{a}_{s},m^{b}_{s})$ in the absence (a; $K{=}0$) and presence (b; $K{=}0.01$) of nonreciprocal coupling as a function of $J$. Inset of (b): Magnification of the bifurcation diagram around the Hopf bifurcation projected onto $(m^{b}_{s},J)$-plane (see black box). (c) Temporal evolution of the magnetization close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation at $(J,K){=}(0.35991,0.01)$. The magnetization oscillates coherently between $4$ ghost states ${\rm A}{\rightarrow}{\rm B}{\rightarrow}{\rm C}{\rightarrow}{\rm D}$ (for $K{<}0$ the direction is reversed), which eventually terminate in the $4$ respective stable branches at and beyond the SNIPER bifurcation. The life-time $\mathcal{T}$ in these ghost states diverges upon approaching the SNIPER bifurcation at $J_{\rm SNP}(K)$ from the left as $\lim_{J\uparrow J_{\rm SNP}(K)}\mathcal{T}\propto(J_{\rm SNP}(K)-J)^{-1/2}$. (d) Phase diagram of $m^{\mu}_{s}$: the gray region depicts the regime of coherent oscillations, with the black lines indicating isolines with fixed oscillation period, and the blue line indicating the Hopf bifurcations where $q_{s}{=}q_{\rm crit}{\equiv}1/(z{-}1)$; in the light green region the magnetization is stationary and nonzero. In all panels we consider $z=4$.
@@
Non-Linear analysis.—Going beyond linear stability, we perform a non-linear analysis of Eqs. (5) via numerical continuation [63]. The resulting bifurcation diagrams are shown in Fig. 2(a,b) and the complete phase diagram in Fig. 2(d), which we now explain in detail.
@@
We start with the non-interacting case with $K=0$ (see Fig. 2a). For small values of $J$, there exists only one (trivial) stable steady state with $m^{\mu}_{s}=0$, as explained in the previous section. Increasing $J$ to $\ln{(z/(z-2))}/2$ we find a pitchfork bifurcation (red dot in Fig. 2a), which coincides with $q_{s}(J,0)=q_{\rm crit}$, and beyond which the trivial stable state becomes unstable. At the pitchfork bifurcation, $4$ stable branches (blue lines in Fig. 2a) and $4$ unstable branches (red dashed lines in Fig. 2a) emerge. Unstable branches have zero magnetization in one of the lattices, while stable branches exhibit nonzero equilibrium magnetization due to spontaneously broken symmetry in both lattices.
@@
Upon setting $K\neq 0$, the pitchfork bifurcation turns into a Hopf bifurcation (blue dot in Fig. 2b), whose $J$-value depends on $K$ through the relation $q_{s}(J,K)=q_{\rm crit}$ (blue line in Fig. 2d; see [55] for explicit results). Increasing $J$ beyond the Hopf bifurcation, there is a regime with coherent oscillations (gray area in Fig. 2b,d). In Fig. 2d we identify the isolines of fixed period of the oscillations (black lines). Upon further increasing $J$ at fixed $K$, we observe a non-linear transition from coherent oscillations to a non-zero stationary magnetization (light green region in Fig. 2d), which is set by a degenerate saddle-node-infinite-period (SNIPER) bifurcation (green dots in Fig. 2b). Approaching the SNIPER bifurcation from below, the magnetization in both lattices start to oscillate between four long-lived ghost states (see Fig. 2c). These $4$ long-lived states correspond to the virtual configuration, in which the respective lattices exert a “static” magnetic field on each other, and are characterized by a critical slowing down of the dynamics in the vicinity of the impending SNIPER bifurcation. The residence time within these ghost states diverges as (see [55])
@@
where $J_{\rm SNP}(K)$ is the $J$-value of the SNIPER bifurcation at a given $K$ (green line in Fig. 2d). At the SNIPER bifurcation, one unstable and stable branch emerge, resulting in $4$ unstable branches (red dashed lines in Fig. 2b) and $4$ stable branches (blue lines in Fig. 2b). Note that these stable states with non-zero stationary magnetization exist also in the finite square lattice system (see proof in [55]).
@@
The bifurcation diagram obtained with the mean-field approximation has similar qualitative features as in Fig. 2a,b (i.e, a Hopf and SNIPER bifurcation), however, the phase diagram displays a constant Hopf line at a fixed $J$, independent of $K$ (see [48] and Fig. A1). In Fig. 2d we see that the Hopf line with the BG approximation is $K$-dependent, closely resembling the empirical phase diagram on the qubic lattice (see [48]).
@@
Spatio-temporal dynamics.—The results in Fig. 1d-f reveal a high degree of local order in the coherent oscillatory regime. To systematically analyze the spatio-temporal patterns in states with coherent oscillations, we performed discrete-time Monte-Carlo simulations of the nonreciprocal Ising system with $2\times N\approx 3\times 10^{3}$ spins on the all-to-all ($z{=}N$), the $z{=}4$ Bethe, and the square ($z{=}4$) lattices with periodic boundary conditions. The respective systems are described in detailed in [55]. For a consistent notion of “spatial scale” on all lattices, we perform a graph-spectral analysis [64].
Figure 3: (a-c) Spectral density $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ of modes of coherent oscillations (Eq. (10)) on the (a) mean-field, (b) Bethe, and (c) square lattice (see [55] for videos of the dynamics on the lattices). Note that the nonzero eigenvalues of the mean-field lattice are degenerate, which explains the white region in (a). The values for $J$ are chosen in the respective oscillatory regimes (see [55] for details), and $K=0.3$ for all three lattices. $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ is averaged over $500$ independent trajectories, and $\omega_{0}={\rm argmax}\langle|\hat{\Psi}^{a}_{1}(\omega)|\rangle$ is the natural oscillation frequency. Resonances are visible at multiplies of $2\omega_{0}$, since we ignore the sign of projections. Insets: Temporal development of the projected microscopic states onto the eigenvectors of the Laplacian matrix of respective lattices. Colors denote three selected eigenvalues (and thus spatial scales) indicated by the dots in the main plot.
@@
Let $\mathbf{L}$ be the $N{\times}N$ symmetric Laplacian matrix of one of the above graphs with elements $L_{ii}{=}z$, and $L_{ij}{=}-1$ when spin $i$ and $j$ are connected and $L_{ij}{=}0$ otherwise. The Laplacian has $N$ orthonormal eigenvectors $\mathbf{L}\boldsymbol{\psi}_{k}{=}l_{k}\boldsymbol{\psi}_{k},\ k\in\{1,...,N\}$ with corresponding eigenvalues $l_{k}$ ordered as $l_{1}{\leq}l_{2}{\leq}\ldots{\leq}l_{N}$. The lowest eigenvalue, corresponding to $\boldsymbol{\psi}_{1}{=}N^{-1/2}(1,\ldots,1)^{\rm T}$, vanishes, i.e., $l_{1}{=}0$ [65]. For the mean-field lattice, all remaining $N{-}1$ eigenvalues are degenerate, $l_{2}=\ldots=l_{N}=N$ [66], but not for the square and Bethe lattices (see [55]).
@@
We express the microscopic state of the lattice $\mu$ at time-step $n\in\{0,...,n_{\rm max}\}$ as a column vector $\boldsymbol{\sigma}^{\mu}(n)=(\sigma^{\mu}_{1}(n),...,\sigma^{\mu}_{N}(n))^{T}$ and project it onto the respective eigenvectors, $\Psi^{\mu}_{k}(n)\equiv\boldsymbol{\psi}_{k}^{T}\boldsymbol{\sigma}^{\mu}(n)$. These spatial modes are shown as insets in Fig. 3a-c, where one can already see that for the Bethe and square lattice oscillations are pronounced on large scales (i.e. for small $k$) and severely suppressed on small scales (large $k$). In the mean-field system, the magnitude of (stochastic) oscillations is equal on all scales (all $k$) since the spectrum of $\mathbf{L}$ is trivial.
@@
To unravel the spatio-temporal structure, we compute the spectral density via the discrete Fourier transform,

### v3

@@ III Linear analysis and the Hopf bifurcation
We focus on the symmetric nonreciprocal setting $J_{a}=J_{b}=J$ and $K_{a}=-K_{b}=K$, also known as the perfectly nonreciprocal Hanai (2024) setting, while keeping $z$ general. We start with the linear stability analysis of Eqs. (8)-(10) around their steady state. The trivial steady state is given by $m^{\mu}_{s}=0$, $q^{ab}_{s}=0$, and
$q^{\mu\mu}_{s}\equiv q_{s}(J,K),$ (16)
which is explicitly given in Appendix C for various values of $z$. Up to first order, small perturbations $\delta\mathbf{m}(t)\equiv(\delta m^{a}(t),\delta m^{b}(t))$ decouple from perturbations $\delta\mathbf{q}(t)\equiv(\delta q^{aa}(t),\delta q^{bb}(t),\delta q^{ab}(t))$, and we obtain the linear equation
$\tau\frac{{\rm d}\delta\mathbf{m}(t)}{{\rm d}t}{=}\begin{pmatrix}M_{1}(q_{s};J,K)&-M_{2}(q_{s};J,K)\\ M_{2}(q_{s};J,K)&M_{1}(q_{s};J,K)\end{pmatrix}\delta\mathbf{m}(t).$ (17)
The linear equation for $\delta\mathbf{q}(t)$ is given in  [^2] See Supplementary Material at […] and does not play any further role here up to linear order. The elements of the linear stability matrix read
$\!\!M_{1}(q_{s};J,K)$ $=\frac{q_{s}/q_{\rm crit}-1}{1+q_{s}}[1-2\sum\nolimits_{l,n}\!\!\overline{\mathcal{P}}^{+}_{l}\tanh(U^{a}_{l,n})],$
$\!\!M_{2}(q_{s};J,K)$ $=\sum\nolimits_{l,n}\!(2n{-}1)[\overline{\mathcal{P}}^{+}_{l}{+}\overline{\mathcal{P}}^{-}_{l}]\tanh(U^{a}_{l,n}),$ (18)
where
$\overline{\mathcal{P}}^{\pm}_{l}(q_{s})\equiv\mathcal{C}^{z}_{l}(1\mp q_{s})^{z-l}(1\pm q_{s})^{l}$ (19)
are the probabilities (14) evaluated at steady-state values, and we introduced the critical local order
$q_{\rm crit}\equiv\frac{1}{z-1},$ (20)
@@
which only depends on the coordination number of the lattice, and sets a critical value for the steady-state local order. The solution of the linear stability equation can be written as
$\mathbf{\delta m}(t)=\sum_{k=\pm}\mathcal{A}_{k}\mathrm{e}^{\lambda_{k}t/\tau}\boldsymbol{\nu}_{k},$ (21)
where $\mathcal{A}_{\pm}$ are set by the initial conditions, $\boldsymbol{\nu}_{\pm}=(\mp{\rm i},1)^{\rm T}$ are the eigenvectors of the linear stability matrix, and $\lambda_{\pm}(q_{s};J,K)$ the corresponding eigenvalues
$\lambda_{\pm}(q_{s};J,K)=M_{1}(q_{s};J,K)\pm{\rm i}M_{2}(q_{s};J,K),$ (22)
@@
$\rm i$ being the imaginary unit. Since ${M_{2}(q_{s};J,K\neq 0)\neq 0}$ (see proof in Appendix D), the eigenvalues are complex for $K\neq 0$ (see dashed lines in Fig. 1b), resulting in oscillatory perturbations. The Hopf bifurcation Strogatz (2018), also called type-${\rm II}_{\rm o}$ instability Cross and Hohenberg (1993), occurs when complex conjugate eigenvalues transit the imaginary axis in the complex plane. According to Eq. (22) this occurs when $M_{1}(q_{s};J,K)=0$ implying ${q_{s}(J,K)=q_{\rm crit}}$ as seen from Eq. (18). The Hopf bifurcation is thus set by the critical value $q_{\rm crit}$ for local order, and for $q_{s}>q_{\rm crit}$ we have ${\rm Re}(\lambda_{\pm})>0$ (solid line in Fig. 1b). In other words, when spins on the respective lattices are sufficiently aligned, a transition to an oscillatory state occurs as shown in Fig. 1c,d.
The critical value of the local order that determines the onset of coherent oscillations is our second main result that generalizes to other approximation schemes beyond the mean-field approximation (see Appendix F). After sufficient local order is attained within the lattices, the frustration due to the nonreciprocal coupling gives rise to coherent oscillations: for $K>0$ a spin $\sigma^{a}_{i}$ wants to align with $\sigma^{b}_{i}$ that, in turn, tends to misalign with $\sigma^{a}_{i}$. This dynamical frustration results in an oscillatory motion of the order parameters Hanai (2024). Notably, for the one-dimensional lattice ($z=2$) we find $q_{\rm crit}=1$, which, in contrast to the mean-field prediction Avni et al. (2023) (see also Appendix E), correctly implies the nonexistence of a Hopf bifurcation.
@@
Figure 2: (a,b) Bifurcation diagram of the magnetization $(m^{a}_{s},m^{b}_{s})$ in the absence (a; $K{=}0$) and presence (b; $K{=}0.01$) of nonreciprocal coupling as a function of $J$. Inset of (b): Magnification of the bifurcation diagram around the Hopf bifurcation projected onto the $(m^{b}_{s},J)$-plane (see black box). (c) Temporal evolution of the magnetization close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation at $(J,K){=}(0.35991,0.01)$. The magnetization oscillates coherently between $4$ ghost states ${\rm A}{\rightarrow}{\rm B}{\rightarrow}{\rm C}{\rightarrow}{\rm D}$ (for $K{<}0$ the direction is reversed), which eventually terminate in the $4$ respective stable branches at and beyond the SNIPER bifurcation. The life-time $\mathcal{T}$ in these ghost states diverges upon approaching the SNIPER bifurcation at $J_{\rm SNP}(K)$ (see Fig. 3). (d) Phase diagram of $m^{\mu}_{s}$: the gray region depicts the regime of coherent oscillations, with the black lines indicating isolines with fixed oscillation period, and the blue line indicating the Hopf bifurcations where $q_{s}=q_{\rm crit}$ [Eq. (20)]; in the light green region the magnetization is stationary and nonzero. In all panels we consider $z=4$.
IV Nonlinear analysis and the SNIPER bifurcation
@@
Going beyond linear stability, we perform a nonlinear analysis of Eqs. (8)-(10) through numerical continuation Dhooge et al. (2003). The resulting bifurcation diagrams are shown in Fig. 2(a,b) and the complete phase diagram in Fig. 2(d), which we now explain in detail.
@@
We start with the non-interacting case with $K=0$ (see Fig. 2a). For small values of $J$, there exists only one (trivial) stable steady state with $m^{\mu}_{s}=0$, as explained in the previous section. Increasing $J$ to $\ln{(z/(z-2))}/2$ we find a pitchfork bifurcation (red dot in Fig. 2a), which coincides with $q_{s}(J,0)=q_{\rm crit}$, and beyond which the trivial stable state becomes unstable. At the pitchfork bifurcation, $4$ stable branches (blue lines in Fig. 2a) and $4$ unstable branches (red dashed lines in Fig. 2a) emerge. Unstable branches have zero magnetization in one of the lattices, while stable branches exhibit nonzero equilibrium magnetization because of spontaneously broken symmetry in both lattices.
Figure 3: Algebraic divergence of the residence time $\mathcal{T}$ within a ghost state close to the degenerate saddle-node-infinite-period (SNIPER) bifurcation. Approaching the SNIPER bifurcation from below, the residence time diverges algebraically according to Eq. (23). Here, we have $K=0.1$ and $z=4$, for which the SNIPER bifurcation occurs at $J_{\rm SNP}\approx 0.4477$. Results are obtained with the continuation package MATCONT Dhooge et al. (2003).
Figure 4: (a-c) Spectral density $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ of modes of coherent oscillations [Eq. (24)] on the (a) mean-field, (b) Bethe, and (c) square lattice (see Note (2) for animations of the dynamics on the lattices). Note that the nonzero eigenvalues of the mean-field lattice are degenerate, which explains the white region in (a). The values for $J$ are chosen in the respective oscillatory regimes (see Appendix G for details), and $K=0.3$ for all three lattices. $\langle|\hat{\Psi}^{a}_{k}(\omega)|\rangle$ is averaged over $500$ independent trajectories, and $\omega_{0}={\rm argmax}\langle|\hat{\Psi}^{a}_{1}(\omega)|\rangle$ is the natural oscillation frequency. Resonances are visible at multiplies of $2\omega_{0}$, since we ignore the sign of projections. Insets: Temporal development of the projected microscopic states onto the eigenvectors of the Laplacian matrix of respective lattices. Colors denote three selected eigenvalues (and thus spatial scales) indicated by the dots in the main plot.
@@
Upon setting $K\neq 0$, the pitchfork bifurcation turns into a Hopf bifurcation (blue dot in Fig. 2b), whose $J$-value depends on $K$ through the relation $q_{s}(J,K)=q_{\rm crit}$ (blue line in Fig. 2d; see Appendix C for explicit results). Increasing $J$ beyond the Hopf bifurcation, there is a regime with coherent oscillations (gray area in Fig. 2b,d). In Fig. 2d we identify the isolines of fixed period of the oscillations (black lines). Upon further increasing $J$ at fixed $K$, we observe a nonlinear transition from coherent oscillations to a nonzero stationary magnetization (light green region in Fig. 2d), which is set by a degenerate saddle-node-infinite-period (SNIPER) bifurcation (green dots in Fig. 2b). Approaching the SNIPER bifurcation from below, the magnetization in both lattices starts to oscillate between four long-lived ghost states (see Fig. 2c). These $4$ long-lived states correspond to the virtual configuration, in which the respective lattices exert a quasi-static magnetic field on each other, and are characterized by a critical slowing down of the dynamics in the vicinity of the impending SNIPER bifurcation. The residence time within these ghost states diverges algebraically as (see Fig. 3)
@@
where $J_{\rm SNP}(K)$ is the $J$-value of the SNIPER bifurcation at a given $K$ (green line in Fig. 2d). At the SNIPER bifurcation, one unstable and stable branch emerges, resulting in $4$ unstable branches (red dashed lines in Fig. 2b) and $4$ stable branches (blue lines in Fig. 2b). Note that these stable states with nonzero stationary magnetization also exist on the finite square lattice system (see proof in Note (2)).
@@
The bifurcation diagram obtained with the mean-field approximation has similar qualitative features as in Fig. 2a,b (i.e, a Hopf and SNIPER bifurcation), however, the phase diagram displays a constant Hopf line at a fixed $J$, independent of $K$ (see Avni et al. (2023) and Fig. E1). In Fig. 2d we see that the Hopf line with the BG approximation is $K$ dependent, closely resembling the empirical phase diagram on the qubic lattice (see Avni et al. (2023)).
V Spatiotemporal dynamics
@@
The results in Fig. 1d-f reveal a high degree of local order in the coherent oscillatory regime. To systematically analyze spatiotemporal patterns in states with coherent oscillations, we performed discrete-time Monte Carlo simulations of the nonreciprocal Ising system with $2\times N\approx 3\times 10^{3}$ spins on the all-to-all ($z{=}N$), the $z{=}4$ Bethe, and the square lattice with periodic boundary conditions. The respective systems are described in detail in Appendix G. For a consistent notion of “spatial scale” on all lattices, we perform a graph-spectral analysis van der Kolk et al. (2023).
@@
Let $\mathbf{L}$ be the $N{\times}N$ symmetric Laplacian matrix of one of the above graphs with elements $L_{ii}{=}z$, and $L_{ij}{=}-1$ when spin $i$ and $j$ are connected and $L_{ij}{=}0$ otherwise. The Laplacian has $N$ orthonormal eigenvectors $\mathbf{L}\boldsymbol{\psi}_{k}{=}l_{k}\boldsymbol{\psi}_{k},\ k\in\{1,...,N\}$ with corresponding eigenvalues $l_{k}$ ordered as $l_{1}{\leq}\ldots{\leq}l_{N}$. The lowest eigenvalue, corresponding to $\boldsymbol{\psi}_{1}{=}N^{-1/2}(1,\ldots,1)^{\rm T}$, vanishes, i.e., $l_{1}{=}0$ Anderson Jr and Morley (1985). For the mean-field lattice, all remaining $N{-}1$ eigenvalues are degenerate, $l_{2}{=}\ldots{=}l_{N}{=}N$ Cohen and Havlin (2010), but not for the square and Bethe lattices (see Fig. G1).
@@
We express the microscopic state of the lattice $\mu$ in timestep $n\in\{0,...,n_{\rm max}\}$ as a column vector $\boldsymbol{\sigma}^{\mu}(n)=(\sigma^{\mu}_{1}(n),...,\sigma^{\mu}_{N}(n))^{\rm T}$ and project it onto the respective eigenvectors, $\Psi^{\mu}_{k}(n)\equiv\boldsymbol{\psi}_{k}^{\rm T}\boldsymbol{\sigma}^{\mu}(n)$. These spatial modes are shown as insets in Figs. 4a-c, where we see that for the Bethe and square lattice oscillations are pronounced on large scales (small $k$) and suppressed on small scales (large $k$). In the mean-field system the spatial modes are equal on all scales due to the degenerate eigenvalues.
@@
To unravel the spatiotemporal structure, we compute the spectral density via the discrete Fourier transform,